Solve the system of linear equations using any method.
2x + y = 5 and 4x + 3y = 11

Answer by mananth(16949)

2x + y = 5 .................1
4x + 3y = 11................2
...
multiply equation 1 by -2
we get
-4x-2y=-10
add to equation2
3y-2y=11-10
y = 1
..
plug value of y in equation 1
2x+1=5
2x=5-1
2x=4
x=2
..
m.ananth@hotmail.ca